Interest-only mortgage with monthly payments and loan amount: $56,000; Term: 15 years; Annual interest rate: 7.5%

  1. What is the total payment in the 180th month? [7 points]
  2. What is the outstanding balance at the end of 10 years (120 months)? [7 points]
  3. What is the total interest payment during the entire loan term (180 months)? [6 points]
Answer #1

As this is interest only loan payment will only be interest for 180 months and final loan amount after 180 months

1.
=56000*7.5%/12=350

2.
=56000=56000

3.
=56000*7.5%*15=63000